OpenStudy (anonymous):

What is the coordinates to the graph y =|x +3| + 2

OpenStudy (anonymous):

- |x + 3| + 2

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=y+%3D |x+%2B3|+%2B+2

OpenStudy (anonymous):

the graph of the function extends along all real numbers

OpenStudy (anonymous):

on the left hand side of x=-3, y=-(x+3)+2. On the right side of x=-3, y=(x+3)+2=x+5

OpenStudy (anonymous):

i hope this helps. take care